OpenStudy (anonymous):

Write an equation in point-slope form of a line that passes through the point (-5, 7) and whose slope is 6. y – 7 = 6(x + 5) y – 7 = 6(x – 5)

OpenStudy (anonymous):

Point slope form is \[y-y1=m(x-x1)\] so it would be y-7=6(x-5)